OpenStudy (anonymous):

The area of a square garden is 128 m squared. How long is the diagonal?

OpenStudy (anonymous):

a = d^2 / 2

OpenStudy (amistre64):

area of a square = side*side 128 = s^2 s = sqrt(128)

OpenStudy (anonymous):

128 = d^2 / 2 d^2 = 256 d=16 m

OpenStudy (amistre64):

err... the diag is just what sqrt(128+128) then :)